A Norman window with a perimeter of 20 units has the shape of a rectangle surmounted by a
semi-circle. If its base is x units long and the height of the rectangle is y units long, write a mathematical model for its area in terms of x.
1
Expert's answer
2020-09-14T18:54:35-0400




let P perimeter and S area

r = x/2 semi-circle where r radius

P = x+2*y +"\pi" *x/2

y = (P -"\pi" *x/2-x)/2 = (10 - "\pi" *x/4- x/2)

S = x*y +"\pi" *(x/2)2/2

S = x* (10 - "\pi" *x/4- x/2)/2 + "\pi" *(x/2)2/2 = 10x - "\pi" *x2/8 -x2/4 + "\pi" *x2/8 = 10x - x2/4